i need to solve the system y=4x+12 and y=-2x+6
Answer by stanbon(75887)

i need to solve the system y=4x+12 and y=-2x+6
Since you have 2 expressions for "y", equate them
and solve for "x":
4x+12 = -2x+6
6x = 18
x = 3
---
Now solve for "y":
y = 4x+12
y = 4*3+12
y = 24
--------
Cheers,
stan H.
==============